# SEAL: Steerable Reasoning Calibration of Large Language Models for Free

### v2

@@ 5.2 End-to-End Results
We compare SEAL with another training-free method, namely logits penalty Wang et al. (2025). Following their setup, we reduce the logits of tokens associated with characteristic reflection or transition words, such as ‘wait’ and ‘alternatively.’ The adjustment values were selected from {-1, -3, -10, -20}, with -3 generally yielding the best results. For each task, we report both the average accuracy and token count across the test set to assess the effectiveness and efficiency of SEAL. To ensure reproducibility, we use greedy search during decoding. Results using sampling are additionally reported in the appendix D.
@@
SEAL demonstrates superior performance and significant efficiency gains. As shown in Table 1 and 2, across diverse models and tasks, SEAL not only improves accuracy but also reduces response length, enhancing reasoning efficiency. For example, SEAL enhances performance by up to 14.1% in accuracy while reducing token usage by 28.8% on hard problems in the Math-500 benchmark. This improvement stems from the model’s ability to avoid excessive rechecking and minimize unnecessary transition thoughts after steering, leading to a more direct path to the final answer. More results on sequence length comparisons can be found in Appendix E. Additionally, SEAL prevents the model from becoming trapped in an endless cycle of rechecking and switching, further optimizing the reasoning process.
@@ 5.3.1 Ablation Study about the Steering Type
Table 4 presents the results of steering different reasoning types. We observe that weakening execution thoughts leads to performance degradation, as execution plays a crucial role in the reasoning process. In contrast, reducing the influence of reflection and transition thoughts individually can still enhance performance. Moreover, enabling the steering vector to weaken both reflection and execution yields the most significant improvements. Consequently, the optimal reasoning formulation is given by .
@@ Appendix C SEAL Performance on In-Distribution Steering Vector
Appendix D Results with Sampling-Based Decoding
@@
Appendix E Comparison of Generated Sequence Lengths
@@
Appendix F Thoghts Patten Demonstration Examples

### v3

@@ 5.2 End-to-End Results
We compare SEAL with another training-free method, namely logits penalty Wang et al. (2025). Following their setup, we reduce the logits of tokens associated with characteristic reflection or transition words, such as ‘wait’ and ‘alternatively.’ The adjustment values were selected from {-1, -3, -10, -20}, with -3 generally yielding the best results. For each task, we report both the average accuracy and token count across the test set to assess the effectiveness and efficiency of SEAL. To ensure reproducibility, we use greedy search during decoding. Results using sampling are additionally reported in the appendix E.
@@
SEAL demonstrates superior performance and significant efficiency gains. As shown in Table 1 and 2, across diverse models and tasks, SEAL not only improves accuracy but also reduces response length, enhancing reasoning efficiency. For example, SEAL enhances performance by up to 14.1% in accuracy while reducing token usage by 28.8% on hard problems in the Math-500 benchmark. This improvement stems from the model’s ability to avoid excessive rechecking and minimize unnecessary transition thoughts after steering, leading to a more direct path to the final answer. More results on sequence length comparisons can be found in Appendix F. Additionally, SEAL prevents the model from becoming trapped in an endless cycle of rechecking and switching, further optimizing the reasoning process.
@@ 5.3.1 Ablation Study about the Steering Type
Table 4 presents the results of steering different reasoning types. We observe that weakening execution thoughts leads to performance degradation, as execution plays a crucial role in the reasoning process. In contrast, reducing the influence of reflection and transition thoughts individually can still enhance performance. Moreover, enabling the steering vector to weaken both reflection and transition yields the most significant improvements. Consequently, the optimal reasoning formulation is given by .
@@ 6 Conclusion
Acknowledgment
Z. Wang is in part supported by NSF Awards 2523383 (AIMing), the NSF AI Institute for Foundations of Machine Learning (IFML), and an Intel single PI SRS gift funding.
@@ B.2 Ablation on different criteria
B.3 LLM-Based Labeling and Comparison with Heuristic Method
In addition to the keyword-based categorization introduced above, we also experimented with using a frontier model (GPT-4o) to label thought types. Given the full reasoning context and the current step, we prompted the LLM to classify each thought as execution, reflection, or transition.
Interestingly, our experiments showed that the keyword-based method achieves slightly better end-task performance in SEAL as shown in Table 7.
| Method | Accuracy (%) |
| Baseline | 85.8 |
| SEAL (Keyword-based) | 89.4 |
| SEAL (LLM-based) | 88.6 |
That said, the LLM-based method remains appealing due to its ease of adaptation to new domains, languages, or expanded categories, and could be especially useful in scenarios where heuristic rules are hard to define or scale.
@@ Appendix C SEAL Performance on In-Distribution Steering Vector
Appendix D Generalization to Open-Ended Planning Tasks
To further evaluate the generalizability of SEAL beyond structured domains such as mathematics and programming, we conducted additional experiments on a more open-ended reasoning benchmark: the NaturalPlan dataset( Zheng et al. (2024)), which focuses on natural language calendar planning.
Importantly, we reused the reasoning steering vector extracted from the MATH domain, to assess cross-domain transferability. As shown in Table 9, SEAL achieved consistent improvements in both accuracy and response length compared to the baseline, indicating that the latent patterns of reflection and transition overuse also appear in more open-ended planning contexts.
| Method | Accuracy (%) | Avg. Response Length |
| Baseline | 14.9 | 6341 |
| SEAL (Math Vector) | 18.6 | 3531 |
Appendix E Results with Sampling-Based Decoding
@@
Appendix F Comparison of Generated Sequence Lengths
@@
Appendix G Disentangling SEAL’s Gains from Truncation Effects
While SEAL demonstrates performance improvements, one may question whether the gains stem primarily from suppressing redundant thoughts or simply from shortening the sequence length, which might reduce token-level errors. To disentangle these effects, we conducted a controlled comparison against a baseline that forcibly limits reasoning length without semantic steering.
Specifically, we introduced a fixed “think budget” mechanism, where the model is instructed to insert an </think> token once the total token count in the reasoning process exceeds a preset budget. After this point, the model proceeds directly to the answer summarization phase. We set the budget to 3500 tokens, closely matching the average output length of SEAL to ensure a fair comparison.
| Method | Accuracy (%) | Avg. Response Length |
| Baseline | 85.8 | 3389 |
| Baseline + Think Budget = 3500 | 85.0 | 2540 |
| SEAL | 89.4 | 2661 |
As shown in Table 11, the truncation-based baseline underperforms SEAL, despite producing similarly shorter outputs. This demonstrates that SEAL’s advantage is not merely due to sequence shortening. Instead, it effectively guides the model to suppress specific types of non-productive reasoning, such as unnecessary reflection and transition thoughts, leading to better-calibrated and more accurate reasoning trajectories.
Appendix H Thoghts Patten Demonstration Examples